# Axi-majoron for almost everything

### v2

@@ 4 Leptogenesis
$\dot{\theta}_{e}\sim\theta_{\rm ini}\left(\frac{m_{a,e}^{2}}{H_{e}^{2}}\right)H_{e}\simeq\frac{\sqrt{3}}{3}\theta_{\rm ini}\alpha\xi_{\phi}\lambda_{\phi}^{\rm ref}M_{\rm P}\ ,$ (4.2)
@@
with $\phi_{e}\simeq M_{\rm P}/\sqrt{\xi_{\phi}}$ [40] was used in the last approximation. Also, for convenience, one may define the effective temperature of the radiation-like energy component at the end of inflation as
@@
It seems that, when $\lambda_{\phi}$ and $\xi_{\phi}$ satisfy the relation in Eq. (3.51) to provide inflationary observables matching observations, $Y_{\Phi,e}$ depends only on $\alpha$, modulo the dependence on the initial condition, $\theta_{\rm ini}$. It should be noted that Eq. (4.5) provides just a qualitative idea of the maximal possibility of $Y_{\Phi,e}$, though. The exact parametric dependence of $Y_{\Phi,e}$ is non-trivial, since it involves the motion of $\Phi$ during inflation that is affected by the initial position of $\phi$, i.e., $\phi_{\rm ini}$ in addition to $\theta_{\rm ini}$. For later convenience, we may express $\dot{\theta}_{e}$ in terms of $Y_{\Phi,e}$ as
@@
Fig. 8 shows numerical estimations of $Y_{\Phi}$ well after inflation as functions of $\theta_{\rm ini}$ for various values of $\alpha$.
@@
The very different behaviors of the $\alpha=10^{-2}$ case in the left and right panels of the figure are due the fact that the angular motion during inflation is sizable for a large $\alpha$ so that, if $\phi_{\rm ini}$ is quite large, the field configuration can move the direction of the minimum along angular direction first, and towards the origin, afterward. From the figure, one can see that, for $\alpha\lesssim 10^{-2}$ it is possible to obtain a quite large amount of asymmetry at least for $\theta_{\rm ini}=\mathcal{O}(0.1-1)$ which may be quite natural.
The eventual decays of $\Phi$ field transfer the PQ-asymmetry to the asymmetry of RHN-helicity states via the interaction $y_{N}$ responsible for generating the majorana mass of RHNs. One may assign 2-units of lepton number to the Peccei-Quinn field, and one-unit to the right-handed neutrino $\nu_{R}$. Then, both interactions $y_{N}$ and $y_{\nu}$ respects not only $U(1)_{\rm PQ}$ but also $U(1)_{B-L}$, and the PQ-number asymmetry of $\Phi$ field can be regarded as the lepton number asymmetry as $\Delta n_{L}^{\Phi}=2\Delta n_{\Phi}$. In the broken phase of $U(1)_{\rm PQ}$-symmetry, the non-zero majorana mass-term breaks the lepton number conservation. However, as eigenstates of the equation of motion, helicity states are not affected by the majorana mass-term and the asymmetry between the helicity states of RHNs is not changed, especially, in the CMB frame of the homogeneous and isotropic universe [68]. The Yukawa interaction $y_{\nu}$ violates the helicity conservation, but the effect depends on the speed of RHNs. The branching ratio of the channel, $N_{i}^{+}\to\ell+H$, is $\left(1+\beta_{i}\right)/2$ whereas that of the channel, $N_{i}^{+}\to\overline{\ell}+\overline{H}$, is $\left(1-\beta_{i}\right)/2$ with $\beta_{i}$ being the speed of $N_{i}$. Hence, the helicity asymmetry of $N_{i}$ induced from the decay of $\Phi$-field is only partially transferred to the visible sector by the suppression factor $\beta_{i}$. As a result, if $N_{i}$ is never equilibrated, its decay contributes to the lepton number asymmetry of the visible sector as [68]
@@
The late time asymmetry in the visible sector depends not only on the speeds of RHNs but also on the background temperature of the SM thermal bath when RHNs decay. Depending on parameters causing different situations, there are several possibilities for generating baryon number asymmetry of the universe: Affleck-Dine (AD) baryogenesis [29] through the decay chain such as $\Phi\to\nu_{R}\to\ell$ with RHNs out of equilibrium, spontaneous leptogenesis thanks to the background motion of the axi-majoron field with thermal RHNs [69], and resonant leptogenesis [30] from decays of non-thermal RHNs.
@@ 4.1 Peccei-Quinn Affleck-Dine leptogenesis
We consider the case in which the decay of $N_{i}$ happen in non-relativistic regime and the background temperature is much lower than $m_{N_{i}}$ so that washing-out processes are not active. Looking at Eqs. (3.83), (3.84) and (3.88), we see that this is relevant only for the lightest RHN $N_{1}$ for which $m_{\nu_{1}}$ can be much smaller than $m_{\nu}$, and $N_{1}$ should be responsible for recovering the late time SM thermal bath.
A part of the PQ-number asymmetry can be transferred to the visible sector through the helicity states of $N_{1}$, while the contributions from heavier states are washed to a much smaller value. It is possible for the heavy states to decay when they just become non-relativistic while producing SM particles with a temperature $T\gtrsim m_{N}\gtrsim m_{N_{1}}$. In such a thermal bath, the scattering rate of $N_{1}$ to SM particles when $T\gtrsim m_{N_{1}}$ is expected to be $\Gamma_{1,s}=c_{1,s}y_{\nu_{1}}^{2}T$ with $c_{1,s}=\mathcal{O}(10^{-2})$ [68] being a numerical factor including the effect of SM couplings and the number of allowed channels. The expansion rate at $T=m_{N_{1}}$ in our scenario is given by
@@
$\simeq$ $10^{-3}\times\left(\frac{\xi_{\phi}}{10}\right)^{2/3}\left(\frac{0.1}{{\rm B}_{1}}\right)^{1/3}\left(\frac{10^{4}m_{\nu_{1}}}{m_{\nu}}\right)^{2/3}\ ,$
where $m_{\phi,\times}\equiv\sqrt{3\lambda_{\phi}}\phi_{\times}$. Also, the yield of the helicity asymmetry of $N_{1}$ at the epoch is found as (see Appendix E)
@@
$\simeq$ $8.4\times 10^{-11}\times\left(\frac{0.1}{{\rm B}_{1}}\right)^{4/3}\left(\frac{\xi_{\phi}}{10}\right)^{19/6}\left(\frac{10^{4}m_{\nu_{1}}}{m_{\nu}}\right)^{7/6}\left(\frac{\phi_{0}^{\rm ref}}{\phi_{0}}\right)\left(\frac{Y_{\Phi,e}}{10^{-2}}\right)\ ,$
where we used $c_{\rm AD}=0.1,\ \Gamma^{\phi}\approx\Gamma_{a}^{\phi}$, and $g_{*,\rm d}=106.75$ in the second line. Since axi-majorons produced by misalignment are the only contribution to dark matter and kinetic misalignment does not work in our scenario (see Section 5), it is required to have $\phi_{0}\sim\phi_{0}^{\rm ref}$. Also, the validity of Eq. (4.15) is limited to the case of $\phi_{*}<\phi_{\times}$ which puts a limit $\xi_{\phi}\lesssim 100$ (see Eq. (3.71)). Moreover, the constraint of $\Gamma_{N_{1}}\ll H_{1,\rm eq}$ requires ${\rm B}_{1}\gtrsim\left(m_{\nu_{1}}/m_{\nu}\right)^{1/2}$ (see Eq. (3.83)). Therefore, if $c_{\rm AD}$ is not much smaller than unity, $Y_{\Phi,e}$ should be smaller than unity by a couple of orders of magnitude in order to match observation.
@@
with $H_{*}$ and $H_{1,\rm eq}$ given in Section 3.3. The yield of the helicity asymmetry of $N_{1}$ at the epoch in this case is given by
@@
$\simeq$ $6.2\times 10^{-11}\times\left(\frac{0.1}{{\rm B}_{1}}\right)^{4/3}\left(\frac{\xi_{\phi}}{200}\right)^{7/6}\left(\frac{10^{4}m_{\nu_{1}}}{m_{\nu}}\right)^{7/6}\left(\frac{Y_{\Phi,e}}{10^{-6}}\right)\ .$
@@
Figure 10: Parameter for a baryon number asymmetry (red solid line), matching observations $Y_{B}^{\rm obs}\simeq 9\times 10^{-11}$ [71], from the contribution of the lightest RHN via AD-mechanism as a function of $m_{\nu_{1}}$ and $b_{1}$ for $b_{N}=0.1$, and $\phi_{0}=\phi_{0}^{\rm ref}$. Left: $Y_{B,\rm AD}^{<}$ with $Y_{\Phi,e}=10^{-1}$ in Eq. (4.15) for $\xi_{\phi}=10$ leading to $\phi_{*}/\phi_{\times}=9.1\times 10^{-3}$. Right: $Y_{B,\rm AD}^{>}$ with $Y_{\Phi,e}=10^{-6}$ in Eq. (4.18) for $\xi_{\phi}=200$ leading to $\phi_{*}/\phi_{\times}=3.6$. For a smaller $Y_{\Phi,e}$, the red line is pushed up. The blue shaded region may be excluded due to too much dark radiation. Blue lines corresponds to $\Delta N_{\rm eff}=0.1$(dotted), $0.2$(dashed), and $0.5$(solid). Green lines are for $T_{\rm d}/m_{N_{1}}=0.02$(dotted), $0.05$(dashed), and $0.1$(solid).
Fig. 10 shows the parameter space for $Y_{B,\rm AD}$ matching observations. As shown in the figure, we can obtain the right amount of baryon asymmetry if $m_{\nu_{1}}$ is smaller than $m_{\nu}$ by several orders of magnitude while a sizable extra-radiation can also be obtained from hot axi-majorons (see Section 5).
A remark is in order before closing this subsection. The initial asymmetry stored in $\Phi$ is only partially transferred to the visible sector with the help of RHNs. The remaining part of the asymmetry is kept as the background motion of axi-majoron. Since $U(1)_{\rm PQ}$ symmetry is broken only spontaneously in the matter sector, the amount of the net PQ-number asymmetry in a co-moving volume after inflation should be conserved. This means that the loss of asymmetry in the visible sector should appear as the motion of axi-majoron, but has minor impact on the overall homogeneous motion. Meanwhile, if the background axi-majoron motion is large enough, it might become a large source of baryon number asymmetry via spontaneous leptogenesis as discussed in Refs. [41, 72, 69] for example. We discuss this possibility in the next subsection.
@@ 4.2 Spontaneous leptogenesis via axi-majoron
If RHNs are in thermal bath while they are coupled to the derivative of the axi-majoron field, the idea of spontaneous baryogenesis [42] can be applied. In Ref. [69], it was shown that in the presence of a non-zero $\dot{\theta}$, the derivative couplings of the majoron(the axi-majoron in our scenario) to fermion currents can act as a source term for generating non-zero lepton number asymmetry via inverse decay processes ($\ell+H\to N_{i}^{\pm}$) in equilibrium through the Yukawa interaction $y_{\nu}$. In our scenario, due to its nature of sharing the properties of axion and majoron, axi-majoron contributes to QCD sphaleron [73] as well as electroweak sphaleron processes, and Boltzmann equations are to be modified correspondingly. However, once $L$-number violation is frozen when the inverse decay to $N_{i}^{\pm}$ is closed, $B-L$ is also frozen. Also, $B+L$ is eventually driven to cancel axi-majoron contribution in the electroweak sphaleron contribution in Boltzmann equations, and $\dot{\theta}/T$ at the freeze-out epoch of the sphalerons is negligible in our scenario as will be clear in the subsequent discussion. Hence, the eventual baryon and lepton number asymmetries are determined by the non-zero $B-L$ number which is determined in a way similar to Ref. [69]. The effect of QCD/electroweak sphalerons on the results is minor, and we will ignore such a difference since it is easily compensated by other parameters.
@@
$T_{\gamma,{\rm fo}}=\frac{m_{N}}{x_{\rm fo}}=\frac{\sqrt{b_{N}\lambda_{\phi}}}{x_{\rm fo}}\phi_{0}\simeq 2.6\times 10^{6}{\rm GeV}\times b_{N}^{1/2}\xi_{\phi}\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)\ ,$ (4.22)
@@
Hence, $\dot{\theta}T_{\gamma}^{2}\propto a^{-5}$ for ${\tilde{\phi}}_{*}\ll\phi_{0}$. The standard thermal bath is expected to be established by the decay of the lightest RHN $N_{1}$ well before the EWPT. Since $N_{1}$s should dominate the universe for a while as non-relativistic particles, there would be a dilution of pre-existing baryon and lepton number asymmetries. The late time baryon number asymmetry through the spontaneous baryogenesis mechanism is given by
@@
$\frac{{\tilde{H}}_{\rm fo}}{{\tilde{H}}_{1,\rm eq}}$ $=$ $\frac{3c_{\rm inv}}{2}\frac{b_{N}}{b_{1}{\rm B}_{1}^{2}\lambda_{\phi}^{1/2}}\left(\frac{m_{\nu_{N}}\phi_{0}}{v_{h}^{2}}\right)$ (4.26)
$\simeq$ $1.4\times\left(\frac{b_{N}}{b_{1}}\right)\left(\frac{0.1}{{\rm B}_{1}}\right)^{2}\left(\frac{10}{\xi_{\phi}}\right)\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)\ .$
Figure 11: Parameter space of ${\tilde{H}}_{\rm fo}/{\tilde{H}}_{1,\rm eq}=1$(orange dashed line), $\Delta N_{\rm eff}=0.5$(blue solid line), $0.2$(blue dashed line), $0.1$(blue dotted line), and $\phi_{*}/\phi_{\times}=1$(purple dashed line) for $b_{1}=0.05,\ b_{N}=0.1$, and $m_{\nu_{N}}=m_{\nu}=10^{4}m_{\nu_{1}}$. The black dotted line is the upper bound of $\xi_{\phi}$ given in Eq. (3.70). The position of purple line does not depend on $b_{i}$s and $m_{\nu_{i}}$s. A larger/smaller $b_{1}$ with other parameters fixed shifts the orange dashed line to the left/right (see Eq. (4.26)). The blue shaded regions and blue lines are move in the same way, but a smaller $m_{\nu_{1}}$ pushes them to the left.
Fig. 11 shows the parameter spaces for ${\tilde{H}}_{\rm fo}/{\tilde{H}}_{1,\rm eq}$. In the region of ${\tilde{H}}_{\rm fo}\geq{\tilde{H}}_{1,\rm eq}$, one finds
@@
$Y_{B,\rm sb}^{<}$$\approx$ $\frac{3.3c_{B-L}c_{\rm inv}^{2}x_{\rm fo}^{2}}{\pi^{9/2}g_{*,\rm d}^{1/4}g_{*,\rm fo}}\frac{b_{N}^{2}\lambda_{\phi}^{5/4}}{b_{1}}\left(\frac{m_{\nu_{1}}M_{\rm P}}{v_{h}^{2}}\right)^{1/2}\left(\frac{m_{\nu_{N}}M_{\rm P}}{v_{h}^{2}}\right)^{2}Y_{\Phi,e}$ (4.30)
$\simeq$ $1.5\times 10^{-10}\left(\frac{0.05}{b_{1}}\right)\left(\frac{b_{N}}{0.1}\right)^{2}\left(\frac{\xi_{\phi}}{50}\right)^{5/2}\left(\frac{10^{4}m_{\nu_{1}}}{m_{\nu}}\right)^{1/2}\left(\frac{Y_{\Phi,e}}{0.1}\right)\ ,$
@@
$\frac{Y_{B,\rm sb}^{<}}{Y_{B,\rm AD}^{<}}\simeq 55\times{\rm B}_{1}^{1/3}b_{3}{\rm B}_{3}\left(\frac{m_{\nu}}{10^{8}m_{\nu_{1}}}\right)^{2/3}\ .$ (4.31)
It implies that in order for the contribution from spontaneous leptogenesis to become relevant, $m_{\nu_{1}}$ should be smaller than $m_{\nu}$ by many orders of magnitude. However, $Y_{\Phi,e}\lesssim\mathcal{O}(1)$ as shown in Eq. (4.5) and Fig. 8, and hence $Y_{B,\rm sb}^{<}$ is always subdominant, relative to $Y_{B,\rm AD}^{<}$ for a given set of parameters.
@@
That is, the contribution to the baryon number asymmetry in the visible sector is always sub-dominant than the one in Eq. (4.30). Hence, the contribution from AD-mechanism is expected to be always dominant in this case, too, and we do not consider the details of this case further.
As the last possibility arising in spontaneous leptogenesis in our scenario, as shown in Fig. 11, and inferred from Eq. (4.26), if $\xi_{\phi}\gtrsim\mathcal{O}(10^{2})$ with $b_{1}$ not much smaller than $b_{N}$, the case of $H_{\rm fo}<H_{1,\rm eq}$ can arise with $\phi_{*}>\phi_{\times}$, too. The $B-L$ asymmetry at the epoch of $N_{1}$’s decay in this case is given by

### v3

@@ 4 Leptogenesis
$\dot{\theta}_{e}\sim\theta_{\rm ini}\left(\frac{m_{a,e}^{2}}{H_{e}^{2}}\right)H_{e}\simeq\frac{\sqrt{3}}{3}\theta_{\rm ini}\alpha\xi_{\phi,0}\lambda_{\phi}^{\rm ref}M_{\rm P}\ ,$ (4.2)
@@
with $\phi_{e}\simeq M_{\rm P}/\sqrt{\xi_{\phi,0}}$ [45] was used in the last approximation. Also, for convenience, one may define the effective temperature of the radiationlike energy component at the end of inflation as
@@
It seems that, when $\lambda_{\phi}$ and $\xi_{\phi,0}$ satisfy the relation in Eq. (3.58) to provide inflationary observables matching observations, $Y_{\Phi,e}$ depends only on $\alpha$, modulo the dependence on the initial condition, $\theta_{\rm ini}$. It should be noted that Eq. (4.5) provides just a qualitative idea of the maximal possibility of $Y_{\Phi,e}$, though. The exact parametric dependence of $Y_{\Phi,e}$ is nontrivial, since it involves the motion of $\Phi$ during inflation that is affected by the initial position of $\phi$, i.e., $\phi_{\rm ini}$ in addition to $\theta_{\rm ini}$. For later convenience, we may express $\dot{\theta}_{e}$ in terms of $Y_{\Phi,e}$ as
@@
Figure 8 shows numerical estimations of $Y_{\Phi}$ well after inflation as functions of $\theta_{\rm ini}$ for various values of $\alpha$.
@@
The very different behaviors of the $\alpha=10^{-2}$ case in the left and right panels of the figure are due the fact that the angular motion during inflation is sizable for a large $\alpha$ so that, if $\phi_{\rm ini}$ is quite large, the field configuration can move the direction of the minimum along angular direction first, and toward the origin, afterward. From the figure, one can see that, for $\alpha\lesssim 10^{-2}$ it is possible to obtain a quite large amount of asymmetry at least for $\theta_{\rm ini}=\mathcal{O}(0.1-1)$ which may be quite natural.
The eventual decays of $\Phi$ field transfer the PQ-asymmetry to the asymmetry of RHN-helicity states via the interaction $y_{N}$ responsible for generating the majorana mass of RHNs. One may assign 2-units of lepton number to the Peccei-Quinn field, and one-unit to the right-handed neutrino $\nu_{R}$. Then, both interactions $y_{N}$ and $y_{\nu}$ respects not only $U(1)_{\rm PQ}$ but also $U(1)_{B-L}$, and the PQ-number asymmetry of $\Phi$ field can be regarded as the lepton number asymmetry as $\Delta n_{L}^{\Phi}=2\Delta n_{\Phi}$. In the broken phase of $U(1)_{\rm PQ}$-symmetry, the nonzero majorana mass-term breaks the lepton number conservation. However, as eigenstates of the equation of motion, helicity states are not affected by the majorana mass-term and the asymmetry between the helicity states of RHNs is not changed, especially, in the CMB frame of the homogeneous and isotropic universe [73]. The Yukawa interaction $y_{\nu}$ violates the helicity conservation, but the effect depends on the speed of RHNs. The branching ratio of the channel, $N_{i}^{+}\to\ell+H$, is $\left(1+\beta_{i}\right)/2$ whereas that of the channel, $N_{i}^{+}\to\overline{\ell}+\overline{H}$, is $\left(1-\beta_{i}\right)/2$ with $\beta_{i}$ being the speed of $N_{i}$. Hence, the helicity asymmetry of $N_{i}$ induced from the decay of $\Phi$-field is only partially transferred to the visible sector by the suppression factor $\beta_{i}$. As a result, if $N_{i}$ is never equilibrated, its decay contributes to the lepton number asymmetry of the visible sector as [73]
@@
The late time asymmetry in the visible sector depends not only on the speeds of RHNs but also on the background temperature of the SM thermal bath when RHNs decay. Depending on parameters causing different situations, there are several possibilities for generating baryon number asymmetry of the universe: Affleck-Dine (AD) baryogenesis [29] through the decay chain such as $\Phi\to\nu_{R}\to\ell$ with RHNs out of equilibrium, spontaneous leptogenesis thanks to the background motion of the axi-Majoron field with thermal RHNs [74], and resonant leptogenesis [30] from decays of nonthermal RHNs.
@@ 4.1 Peccei-Quinn Affleck-Dine leptogenesis
We consider the case in which the decay of $N_{i}$ happen in nonrelativistic regime and the background temperature is much lower than $m_{N_{i}}$ so that washing-out processes are not active. Looking at Eqs. (3.90), (3.91) and (3.95), we see that this is relevant only for the lightest RHN $N_{1}$ for which $m_{\nu_{1}}$ can be much smaller than $m_{\nu}$, and $N_{1}$ should be responsible for recovering the late time SM thermal bath.
A part of the PQ-number asymmetry can be transferred to the visible sector through the helicity states of $N_{1}$, while the contributions from heavier states are washed to a much smaller value. It is possible for the heavy states to decay when they just become nonrelativistic while producing SM particles with a temperature $T\gtrsim m_{N}\gtrsim m_{N_{1}}$. In such a thermal bath, the scattering rate of $N_{1}$ to SM particles when $T\gtrsim m_{N_{1}}$ is expected to be $\Gamma_{1,s}=c_{1,s}y_{\nu_{1}}^{2}T$ with $c_{1,s}=\mathcal{O}(10^{-2})$ [73] being a numerical factor including the effect of SM couplings and the number of allowed channels. The expansion rate at $T=m_{N_{1}}$ in our scenario is given by
@@
$\simeq$ $10^{-3}\times\left(\frac{\xi_{\phi,0}}{10}\right)^{2/3}\left(\frac{0.1}{{\rm B}_{1}}\right)^{1/3}\left(\frac{10^{4}m_{\nu_{1}}}{m_{\nu}}\right)^{2/3}\ ,$
where $m_{\phi,\times}\equiv\sqrt{3\lambda_{\phi}}\phi_{\times}$. Also, the yield of the helicity asymmetry of $N_{1}$ at the epoch is found as (see Appendix D)
@@
$\simeq$ $8.4\times 10^{-11}\times\left(\frac{0.1}{{\rm B}_{1}}\right)^{4/3}\left(\frac{\xi_{\phi,0}}{10}\right)^{19/6}\left(\frac{10^{4}m_{\nu_{1}}}{m_{\nu}}\right)^{7/6}\left(\frac{\phi_{0}^{\rm ref}}{\phi_{0}}\right)\left(\frac{Y_{\Phi,e}}{10^{-2}}\right)\ ,$
where we used $c_{\rm AD}=0.1,\ \Gamma^{\phi}\approx\Gamma_{a}^{\phi}$, and $g_{*,\rm d}=106.75$ in the second line. Since axi-Majorons produced by misalignment are the only contribution to dark matter and kinetic misalignment does not work in our scenario (see Sec. 5), it is required to have $\phi_{0}\sim\phi_{0}^{\rm ref}$. Also, the validity of Eq. (4.15) is limited to the case of $\phi_{*}<\phi_{\times}$ which puts a limit $\xi_{\phi,0}\lesssim 100$ (see Eq. (3.78)). Moreover, the constraint of $\Gamma_{N_{1}}\ll H_{1,\rm eq}$ requires ${\rm B}_{1}\gtrsim\left(m_{\nu_{1}}/m_{\nu}\right)^{1/2}/\sin\beta$ (see Eq. (3.90)). Therefore, if $c_{\rm AD}$ is not much smaller than unity, $Y_{\Phi,e}$ should be smaller than unity by a couple of orders of magnitude in order to match observation.
@@
with $H_{*}$ and $H_{1,\rm eq}$ given in Sec. 3.3. The yield of the helicity asymmetry of $N_{1}$ at the epoch in this case is given by
@@
$\simeq$ $6.2\times 10^{-11}\times\left(\frac{0.1}{{\rm B}_{1}}\right)^{4/3}\left(\frac{\xi_{\phi,0}}{200}\right)^{7/6}\left(\frac{10^{4}m_{\nu_{1}}}{m_{\nu}}\right)^{7/6}\left(\frac{Y_{\Phi,e}}{10^{-6}}\right)\ .$
@@
Figure 10: Parameter for a baryon number asymmetry (red solid line), matching observations $Y_{B}^{\rm obs}\simeq 9\times 10^{-11}$ [76], from the contribution of the lightest RHN via AD-mechanism as a function of $m_{\nu_{1}}$ and $b_{1}$ for $b_{N}=0.1$, and $\phi_{0}=\phi_{0}^{\rm ref}$. Left: $Y_{B,\rm AD}^{<}$ with $Y_{\Phi,e}=10^{-1}$ in Eq. (4.15) for $\xi_{\phi,0}=10$ leading to $\phi_{*}/\phi_{\times}=9.1\times 10^{-3}$. Right: $Y_{B,\rm AD}^{>}$ with $Y_{\Phi,e}=10^{-6}$ in Eq. (4.18) for $\xi_{\phi,0}=200$ leading to $\phi_{*}/\phi_{\times}=3.6$. For a smaller $Y_{\Phi,e}$, the red line is pushed up. The blue shaded region may be excluded due to too much dark radiation. Blue lines corresponds to $\Delta N_{\rm eff}=0.1$(dotted), $0.2$(dashed), and $0.5$(solid). Green lines are for $T_{\rm d}/m_{N_{1}}=0.02$(dotted), $0.05$(dashed), and $0.1$(solid).
Figure 10 shows the parameter space for $Y_{B,\rm AD}$ matching observations. As shown in the figure, we can obtain the right amount of baryon asymmetry if $m_{\nu_{1}}$ is smaller than $m_{\nu}$ by several orders of magnitude while a sizable extra-radiation can also be obtained from hot axi-Majorons (see Sec. 5).
A remark is in order before closing this subsection. The initial asymmetry stored in $\Phi$ is only partially transferred to the visible sector with the help of RHNs. The remaining part of the asymmetry is kept as the background motion of axi-Majoron. Since $U(1)_{\rm PQ}$ symmetry is broken only spontaneously in the matter sector, the amount of the net PQ-number asymmetry in a comoving volume after inflation should be conserved. This means that the loss of asymmetry in the visible sector should appear as the motion of axi-Majoron, but has minor impact on the overall homogeneous motion. Meanwhile, if the background axi-Majoron motion is large enough, it might become a large source of baryon number asymmetry via spontaneous leptogenesis as discussed in Refs. [46, 77, 74] for example. We discuss this possibility in the next subsection.
@@ 4.2 Spontaneous leptogenesis via axi-Majoron
If RHNs are in thermal bath while they are coupled to the derivative of the axi-Majoron field, the idea of spontaneous baryogenesis [47] can be applied. In Ref. [74], it was shown that in the presence of a nonzero $\dot{\theta}$, the derivative couplings of the Majoron(the axi-Majoron in our scenario) to fermion currents can act as a source term for generating nonzero lepton number asymmetry via inverse decay processes ($\ell+H\to N_{i}^{\pm}$) in equilibrium through the Yukawa interaction $y_{\nu}$. In our scenario, due to its nature of sharing the properties of axion and Majoron, axi-Majoron contributes to QCD sphaleron [78] as well as electroweak sphaleron processes, and Boltzmann equations are to be modified correspondingly. However, once $L$-number violation is frozen when the inverse decay to $N_{i}^{\pm}$ is closed, $B-L$ is also frozen. Also, $B+L$ is eventually driven to cancel axi-Majoron contribution in the electroweak sphaleron contribution in Boltzmann equations, and $\dot{\theta}/T$ at the freeze-out epoch of the sphalerons is negligible in our scenario as will be clear in the subsequent discussion. Hence, the eventual baryon and lepton number asymmetries are determined by the nonzero $B-L$ number which is determined in a way similar to Ref. [74]. The effect of QCD/electroweak sphalerons on the results is minor, and we will ignore such a difference since it is easily compensated by other parameters.
@@
$T_{\gamma,{\rm fo}}=\frac{m_{N}}{x_{\rm fo}}=\frac{\sqrt{b_{N}\lambda_{\phi}}}{x_{\rm fo}}\phi_{0}\simeq 2.6\times 10^{6}{\rm GeV}\times b_{N}^{1/2}\xi_{\phi,0}\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)\ ,$ (4.22)
@@
Hence, $\dot{\theta}T_{\gamma}^{2}\propto a^{-5}$ for ${\tilde{\phi}}_{*}\ll\phi_{0}$. The standard thermal bath is expected to be established by the decay of the lightest RHN $N_{1}$ well before the EWPT. Since $N_{1}$s should dominate the universe for a while as nonrelativistic particles, there would be a dilution of preexisting baryon and lepton number asymmetries. The late time baryon number asymmetry through the spontaneous baryogenesis mechanism is given by
@@
$\frac{{\tilde{H}}_{\rm fo}}{{\tilde{H}}_{1,\rm eq}}$ $=$ $\frac{3c_{\rm inv}}{2}\frac{b_{N}}{b_{1}{\rm B}_{1}^{2}\lambda_{\phi}^{1/2}}\left(\frac{m_{\nu_{N}}\phi_{0}}{v_{u}^{2}}\right)$ (4.26)
$\simeq$ $\frac{1.4}{\sin^{2}\beta}\times\left(\frac{b_{N}}{b_{1}}\right)\left(\frac{0.1}{{\rm B}_{1}}\right)^{2}\left(\frac{10}{\xi_{\phi,0}}\right)\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)\ .$
Figure 11: Parameter space of ${\tilde{H}}_{\rm fo}/{\tilde{H}}_{1,\rm eq}=1$(orange dashed line), $\Delta N_{\rm eff}=0.5$(blue solid line), $0.2$(blue dashed line), $0.1$(blue dotted line), and $\phi_{*}/\phi_{\times}=1$(purple dashed line) for $b_{1}=0.05,\ b_{N}=0.1$, $m_{\nu_{N}}=m_{\nu}=10^{4}m_{\nu_{1}}$, and $\sin\beta=1$ as an example. The black dotted line is the upper bound of $\xi_{\phi,0}$ given in Eq. (3.77). The position of purple line does not depend on $b_{i}$s and $m_{\nu_{i}}$s. A larger/smaller $b_{1}$ with other parameters fixed shifts the orange dashed line to the left/right (see Eq. (4.26)). A smaller $\sin\beta$ shifts the line to the right. The blue shaded regions and blue lines are move in the same way, but a smaller $m_{\nu_{1}}$ pushes them to the left.
Figure 11 shows the parameter spaces for ${\tilde{H}}_{\rm fo}/{\tilde{H}}_{1,\rm eq}$. In the region of ${\tilde{H}}_{\rm fo}\geq{\tilde{H}}_{1,\rm eq}$, one finds
@@
$Y_{B,\rm sb}^{<}$$\approx$ $\frac{3.3c_{B-L}c_{\rm inv}^{2}x_{\rm fo}^{2}}{\pi^{9/2}g_{*,\rm d}^{1/4}g_{*,\rm fo}}\frac{b_{N}^{2}\lambda_{\phi}^{5/4}}{b_{1}}\left(\frac{m_{\nu_{1}}M_{\rm P}}{v_{u}^{2}}\right)^{1/2}\left(\frac{m_{\nu_{N}}M_{\rm P}}{v_{u}^{2}}\right)^{2}Y_{\Phi,e}$ (4.30)
$\simeq$ $\frac{1.5\times 10^{-10}}{\sin^{5}\beta}\left(\frac{0.05}{b_{1}}\right)\left(\frac{b_{N}}{0.1}\right)^{2}\left(\frac{\xi_{\phi,0}}{50}\right)^{5/2}\left(\frac{10^{4}m_{\nu_{1}}}{m_{\nu}}\right)^{1/2}\left(\frac{Y_{\Phi,e}}{0.1}\right)\ ,$
@@
$\frac{Y_{B,\rm sb}^{<}}{Y_{B,\rm AD}^{<}}\simeq\frac{55}{\sin^{5}\beta}\times{\rm B}_{1}^{1/3}b_{3}{\rm B}_{3}\left(\frac{m_{\nu}}{10^{8}m_{\nu_{1}}}\right)^{2/3}\ .$ (4.31)
It implies that, as long as $\beta\gtrsim 1$, in order for the contribution from spontaneous leptogenesis to become relevant, $m_{\nu_{1}}$ should be smaller than $m_{\nu}$ by many orders of magnitude. Meanwhile, $Y_{\Phi,e}\lesssim\mathcal{O}(1)$ as shown in Eq. (4.5) and Fig. 8. Hence, $Y_{B,\rm sb}^{<}$ is always subdominant, relative to $Y_{B,\rm AD}^{<}$ for a given set of parameters.
@@
That is, the contribution to the baryon number asymmetry in the visible sector is always subdominant than the one in Eq. (4.30). Hence, the contribution from AD-mechanism is expected to be always dominant in this case, too, and we do not consider the details of this case further.
As the last possibility arising in spontaneous leptogenesis in our scenario, as shown in Fig. 11, and inferred from Eq. (4.26), if $\xi_{\phi,0}\gtrsim\mathcal{O}(10^{2})$ with $b_{1}$ not much smaller than $b_{N}$, the case of $H_{\rm fo}<H_{1,\rm eq}$ can arise with $\phi_{*}>\phi_{\times}$, too. The $B-L$ asymmetry at the epoch of $N_{1}$’s decay in this case is given by